# medR: Reward Engineering for Clinical Offline Reinforcement Learning via Tri-Drive Potential Functions

### v1

@@ 1 Introduction
Critical care medicine is fundamentally defined by high-stakes, sequential decision-making, RL is a natural fit for optimizing DTRs where an agent must make sequential decisions that adapts to a patient’s changing state over time. Reward serves as the core of RL that not only guides the agent learning but also defines the optimization problem (Zhu et al., ). Small changes in reward function weights can cause profound policy changes(Roggeveen et al., 2024) and bad reward signals affect beyond “win-or-lose” in gaming and can lead to “life-or-death” in healthcare. Errors in sensitive fields such as autonomous driving have already been shown to be problematic (Knox et al., 2023). However, reward engineering in clinical RL is less studied and benchmarked. This underscores an urgent need to design better rewards that align with clinical goals (Basu et al., 2025).
@@
Current rewards in RL for DTRs largely depend on manual and static design. Drawing an analogy to recent advances in LLM, we categorize existing clinical reward mechanisms into three distinct paradigms: sparse Outcome Reward Model (ORM), which assigns scalar rewards based on terminal patient outcomes like survival (Komorowski et al., 2018). The second is dense Process Reward Model (PRM), which depends on dense feedback based on intermediate physiological state changes (Zhang et al., 2024c). To mitigate the sparsity of terminal outcomes, the standard approach augments them with intermediate feedback to form a composite reward ($R_{OPRM}=R_{ORM}+R_{PRM}$) (Raghu et al., 2018; Lu et al., 2024). In this paradigm, clinicians manually select disease-relevant features and encode them using simple heuristic logic. However, the terminal reward is often weighted disproportionately high, overshadowing the dense shaping signals and the selection of specific parameters and weights remains largely ad-hoc, lacking systematic justification (Lu et al., 2024; Zhang et al., 2024c).
@@
Despite efforts in designing various reward signals to accelerate policy learning, there are currently no guidelines on designing effective rewards for clinical RL due to several major challenges. First, outcomes in clinical settings are sparse and delayed. While mortality is the ultimate ground truth outcome, it occurs only at the end of ICU stay which leads to credit assignment problem. Second, the target in a treatment process is often implicit and complicated. It extends beyond binary survival to encompass the quality of recovery. There is a profound gap between the high-dimensional complexity of patient physiology and the scalar reward value required by RL algorithms. This gap leads to misaligned objectives, such as the “Pyrrhic Victory”(Livingston et al., 2019) where an agent optimizes for survival via aggressive interventions or reward hacking. Third, rewards are not verifiable. Unlike robotics or games, clinical RL is strictly offline, precluding online exploration to validate reward signals. Consequently, current rewards rely on brittle, manual heuristics that lack interpretability and fail to generalize across different diseases or datasets.
@@
Recent advancements have identified LLMs as a promising solution to bridge the gap between high-level human intent and low-level reinforcement signals. While early work utilized LLMs as direct evaluators of agent behavior(Kwon et al., 2023), more recent methods focus on using LLMs to generate executable reward function codes(Ma et al., 2023; Xie et al., 2023). We leverage this generative capability to construct a reward function that captures the nuance of DTRs without the prohibitively high cost of manual trial-and-error engineering. Our pipeline leverages LLMs to bridge the semantic gap between medical knowledge and mathematical representation, utilizing a “Tri-Drive” mechanism comprising Survival, Confidence, and Competence to guide both reward generation and offline verification. Our specific contributions are:
@@ 2.1 Intrinsic Motivation
In this work, our reward function design follows the evolutionary perspective proposed in (Singh et al., 2009). They distinguish between reward signals and fitness functions where reward signals are internal scalar values that drive learning within an agent and fitness is the external evaluation that measures an agent’s success. For bounded agents in an offline environment, the optimal reward function is not the fitness function(Singh et al., 2010). The design of reward functions should optimize intrinsic motivations (IM), unlike previous literature about IM that focuses on incentivizing exploration which is dangerous in our offline setting, in our paper, we form it as a threefold drive: the drive to restore homeostasis(survival), the drive to lower uncertainty(confidence), and the drive to minimize cost(competence). In Hull’s Drive Reduction Theory(Hull, 1945), organisms are motivated to maintain physiological equilibrium, we treat homeostasis as a proxy of survival to optimize in the reward function. Unlike curiosity-driven frameworks that encourage exploration under uncertainty(Berlyne, 1960; Schmidhuber, 1991), in order to design a reward function specific to our offline environment(Singh et al., 2009), we conversely suppress uncertainty due to the sensitivity in critical care(Djulbegovic, 2011). White (1959) furthers the learning process of effective interaction with the environment, which is transferrable as effective treatment under clinical setting.
@@ 3.2 Potential Based Reward Function Generation
We design a health potential that reflects the quality of physiological states. Past reward functions often define specific rules that directly scores the action, for example, Reward +1+1 if $70<mbp<80$, which will likely cause reward hacking and encouraging actions that can directly increase blood pressure. Instead, we form it as the gradient of potential to capture the patient state change. To bridge the gap between sparse outcome labels and dense clinical supervision, the reward function $R$ is composed of three distinct drives—Survival, Confidence, and Competence leveraging the intrinsic motivation theory (Singh et al., 2009). We define the reward function as follows:
$\small R(s,a,t,s^{\prime},a^{\prime},t^{\prime})=\left[\gamma\Phi(s_{t+1},t+1)-\Phi(s_{t},t)\right]-\lambda\mathcal{C}(a_{t})$ (1)
$\small\Phi(s_{t},t)=\delta(t)\cdot\sum_{f\in\mathcal{F}}\omega_{f}\cdot\mathcal{S}_{f}(v_{t,f})\cdot\mathcal{U}_{f}(\Delta t_{t,f})$ (2)
@@
We use physiological stabilization as the proxy for survival. By embedding the metadata of critical features as part of the prompt, we personalize the reward design to handle critical patients with comorbidities. The LLM is prompted to design a potential function that rewards states closer to homeostasis. In the offline clinical setting, data irregularity poses a significant risk: a normal heart rate recorded 12 hours ago is not a reliable indicator of current stability. Since we are unable to refresh the data by ordering new lab tests, we utilize the elapsed time $\Delta t$ as indicators of freshness of the data $s_{t}=(o_{t},\Delta t_{t})$. The LLM generates confidence weights for each critical feature that inversely scales the reward based on $\Delta_{t,f}$. This mechanism penalizes stale states, effectively discounting the survival reward when data is uncertain to reduce epistemic uncertainty. We also add a strategic decay to prevent unnecessarily longer ICU stay. To prevent the agent from achieving survival via excessive or harmful interventions, we incorporate a competence component focused on treatment efficiency. Mathematically, this serves as a regularization term: if two treatments result in the same physiological transition $s_{t}\rightarrow s_{t+1}$, the potential function assigns a higher value to the less invasive action, penalizing over-treatment or flailing behavior. By decoupling the accumulating action cost $\mathcal{C}(a)$ from the telescoping potential $\Phi(s,t)$D.1, we effectively formulate the RL objective as the Lagrangian of a Constrained Markov Decision Process (CMDP). The agent maximizes the Lagrangian $\mathcal{L}=\text{Physiology}-\lambda\cdot\text{Toxicity}$, ensuring that high-stability states are only pursued if the cost of the intervention path does not exceed the value of the stabilizationD.2. Instead of manual heuristic tuning, the LLM acts as the logic designer. We feed the definitions of these three drives along with the feature metadata into the LLM, which autonomously synthesizes the logic into executable Python code for the reward function. To mitigate the risk of hallucination and capture diverse clinical perspectives, we generate $N$ distinct candidate functions in parallel, creating a diverse pool of candidates for the subsequent verification phase.
@@ 3.3 Tri-Drive Reward Selection
Since online verification is ethically prohibitive in clinical settings, we cannot evaluate reward candidates by deploying them on patients. Instead, we propose a set of offline fitness metrics to rigorously assess the quality of each generated reward function $R$. We formulate reward design as a Tri-Drive Selection Problem aiming to balance three conflicting objectives based on trajectory-level labeled ground truth: Survival, Confidence, and Competence Fitness.Survival Fitness ($J_{surv}$) validates whether the induced reward signal correlates with actual patient outcomes. We define a ground-truth trajectory score $G(\tau)$ that consists of both mortality and Sequential Organ Failure Assessment(SOFA) score(Lambden et al., 2019) increase compared to baseline(time when admitted to ICU). The fitness is the correlation $\rho$ between the cumulative reward and $G(\tau)$:
@@
$\small h(f)=\begin{cases}\mathbf{1}_{f\in\mathcal{I}}+\mathbf{1}_{f\notin\mathcal{I}}\cdot\sigma\left(k\left[0.5-\frac{d(f,\mathcal{I})}{\text{IQR}}\right]\right)&\text{if }T=N,\\ \sigma\left(k\left[0.5-f\right]\right)&\text{if }T=D_{l},\\ \sigma\left(-k\left[0.5-f\right]\right)&\text{if }T=D_{h},\end{cases}$ (7)
@@
where $\sigma(z)=(1+e^{-z})^{-1}$ is the standard sigmoid function, $d(f,\mathcal{I})$ is the distance from feature $f$ to the interval $\mathcal{I}$, IQR is the interquartile range width used for normalization. $\alpha$ and $k$ are cost parameters, $\bar{a}_{t}$ is the magnitude of the intervention. We employ the Non-Dominated Sorting Genetic Algorithm II (NSGA-II) (Deb et al., 2002) to rank the generated candidates based on these three objectives. Instead of selecting a single weighted average, we identify the Pareto Frontier where no objective can be improved without degrading another. From this frontier, the solution closest to the utopia point representing the best trade-off.
@@ 4.1 Tasks
Sepsis Treatment We utilize the Medical Information Mart for Intensive Care (MIMIC)-IV v3.1(Johnson et al., 2020) database to learn optimal Intravenous (IV) fluid and Vasopressor administration strategies for sepsis patients. We select adult patients (age $\geq$ 18) satisfying the Sepsis-3 criteria (Singer et al., 2016), and exclude patients with missing mortality, demographics, and interventions, extreme missingness (>>60% missing data) or short hospital stays (<<24 hours). We define a septic shock starting from 24 hours before diagnosis and ending at 48 hours post-diagnosis. After filtration, 7,501 patients with 46 features are retained in the dataset. The action space is discrete consisting of two primary interventions: IV fluids and Vasopressors. We discretize IV fluid volume and vasopressor dosage into 5 bins each, resulting in a combined action space of $5\times 5=25$ discrete actions.
@@
Mechanical Ventilation We use the multi-center eICU Collaborative Research Database (eICU-CRD)(Pollard et al., 2018) to optimize ventilator settings for patients with respiratory failure. The cohort includes adult patients who underwent invasive mechanical ventilation (MV) for at least 24 hours. We select patients aged $\geq$16 who received MV for at least 24 hours. We also exclude patients with missing mortality outcomes, demographics, interventions, and those ventilated for >>14 days. The final cohort contains 2168 patients with 41 features. The action space is discrete, controlling three key ventilator parameters: Positive End-Expiratory Pressure (PEEP), Fraction of Inspired Oxygen (FiO2), and Tidal Volume. We define a discrete action set based on clinically relevant adjustments for each parameter, resulting in a composite discrete action space of 18 actions (PEEP: 2 levels, FiO2: 3 levels, Tidal Volume: 3 levels).
@@
Renal Replacement Therapy We apply our method to the AmsterdamUMCdb (Thoral et al., 2021) to optimize continuous renal replacement therapy (RRT) dosing. We select adult patients (age $\geq$ 18) who received CRRT during the ICU stay where the stay is no longer than 60 days. There are 857 patients with 18 features in the cohort. Unlike the previous tasks, the action space here is continuous. The agent controls the effluent flow rate clipped to the range $[0,60]$ ml/kg/h, a critical parameter determining the intensity of dialysis.
@@
For all tasks, clinical time-series data (vitals, labs, medications) are aggregated into 1-hour time steps. Missing values are imputed using forward-filling for time-varying features. For CRRT effluent rate, we imputed values based on RRT sessions, defining a new session whenever the gap between two dialysis records exceeded 8 hours. All continuous state features are normalized to zero mean and unit variance to ensure training stability. For policy training, we use common safe offline RL methods, Batch-Constrained Q-Learning (BCQ)(Fujimoto et al., 2019) for discrete action space tasks (Sepsis and Ventilation), and Implicit Q-Learning (IQL)(Kostrikov et al., 2021) for continuous action space task (RRT).
@@ 4.2.1 Heuristic Reward Design
We implement three standard reward structures commonly used in clinical RL literature. These serve as the ”human-expert” baselines, representing varying degrees of clinical foresight:
Outcome Reward Model (ORM): A purely outcome-oriented function that assigns a non-zero reward only at the terminal step, +100+100 for survival (discharge) and 100-100 for mortality. The value of $\pm 100$ is used in both Sepsis(Komorowski et al., 2018) and Ventilation(Peine et al., 2021) tasks in previous literature. Since there is no past work that utilized outcome reward for RRT task, we adopt the same setting of $\pm 100$ for RRT. This baseline tests the agent’s ability to solve the credit assignment problem without intermediate guidance. We explicitly utilized ICU-discharge mortality rather than 90-day mortality as our terminal state definition in the baselines. We argue that an RL agent can only be held accountable for outcomes within its scope of control. An agent penalized for a death occurring 60 days post-discharge may erroneously learn that stabilizing and discharging a patient is suboptimal, leading to conservative strategies.
Process Reward Model (PRM): A process-oriented function that provides immediate feedback based on step-wise improvements in physiological stability. It incentivizes short-term stabilization without long-term survival awareness. We use the reward function fully based on SOFA score(Zhang et al., 2024c) for Sepsis. For Ventilation, though we note that there is one process reward based on ventilator free days(Yousuf et al., 2025), the reward calculated on our cohort is not dense enough for efficient policy training, we therefore adopt the intermediate reward from the OPRM(Liu et al., 2024). This further indicates that a fixed reward function may not be transferrable between different cohorts even with the same dataset. For RRT, we use the reward function based on Estimated Glomerular Filtration Rate (eGFR)(Zhang et al., 2024a).
Outcome Process Reward Model(OPRM): A weighted combination of the sparse and dense signals. This represents the state-of-the-art manual design, aiming to balance immediate physiological regularization with the ultimate goal of patient survival. We select from current SOTA methods (Perera et al., 2026)for Sepsis and (Liu et al., 2024)for Ventilation. Due to the lack of work on using OPRM for RRT, we choose to combine the PRM(eGFR) with ORM($\pm 100$) as OPRM.

### v2

@@ 1 Introduction
Critical care medicine is fundamentally defined by high-stakes, sequential decision-making, RL is a natural fit for optimizing DTRs where an agent must make sequential decisions that adapts to a patient’s changing state over time. Reward serves as the core of RL that not only guides the agent learning but also defines the optimization problem Zhu et al. . Small changes in reward function weights can cause profound policy changesRoggeveen et al. (2024) and bad reward signals affect beyond “win-or-lose" in gaming and can lead to “life-or-death" in healthcare. Errors in sensitive fields such as autonomous driving have already been shown to be problematic Knox et al. (2023). However, reward engineering in clinical RL is less studied and benchmarked. This underscores an urgent need to design better rewards that align with clinical goals Basu et al. (2025).
@@
Current rewards in RL for DTRs largely depend on manual and static design. Drawing an analogy to recent advances in LLM, we categorize existing clinical reward mechanisms into three distinct paradigms: sparse Outcome Reward Model (ORM), which assigns scalar rewards based on terminal patient outcomes like survival Komorowski et al. (2018). The second is dense Process Reward Model (PRM), which depends on dense feedback based on intermediate physiological state changes Zhang et al. (2024c). To mitigate the sparsity of terminal outcomes, the standard approach augments them with intermediate feedback to form a composite reward ($R_{OPRM}=R_{ORM}+R_{PRM}$) Raghu et al. (2018); Lu et al. (2024). In this paradigm, clinicians manually select disease-relevant features and encode them using simple heuristic logic. However, the terminal reward is often weighted disproportionately high, overshadowing the dense shaping signals and the selection of specific parameters and weights remains largely ad-hoc, lacking systematic justification Lu et al. (2024); Zhang et al. (2024c).
@@
Despite efforts in designing various reward signals to accelerate policy learning, there are currently no guidelines on designing effective rewards for clinical RL due to several major challenges. First, outcomes in clinical settings are sparse and delayed. While mortality is the ultimate ground truth outcome, it occurs only at the end of ICU stay which leads to credit assignment problem. Second, the target in a treatment process is often implicit and complicated. It extends beyond binary survival to encompass the quality of recovery. There is a profound gap between the high-dimensional complexity of patient physiology and the scalar reward value required by RL algorithms. This gap leads to misaligned objectives, such as the “Pyrrhic Victory”Livingston et al. (2019) where an agent optimizes for survival via aggressive interventions or reward hacking. Third, rewards are not verifiable. Unlike robotics or games, clinical RL is strictly offline, precluding online exploration to validate reward signals. Consequently, current rewards rely on brittle, manual heuristics that lack interpretability and fail to generalize across different diseases or datasets.
@@
Recent advancements have identified LLMs as a promising solution to bridge the gap between high-level human intent and low-level reinforcement signals. While early work utilized LLMs as direct evaluators of agent behaviorKwon et al. (2023), more recent methods focus on using LLMs to generate executable reward function codesMa et al. (2023); Xie et al. (2023). We leverage this generative capability to construct a reward function that captures the nuance of DTRs without the prohibitively high cost of manual trial-and-error engineering. Our pipeline leverages LLMs to bridge the semantic gap between medical knowledge and mathematical representation, utilizing a “Tri-Drive” mechanism comprising Survival, Confidence, and Competence to guide both reward generation and offline verification. Our specific contributions are:
@@ 2.1 Intrinsic Motivation
In this work, our reward function design follows the evolutionary perspective proposed in Singh et al. (2009). They distinguish between reward signals and fitness functions where reward signals are internal scalar values that drive learning within an agent and fitness is the external evaluation that measures an agent’s success. For bounded agents in an offline environment, the optimal reward function is not the fitness functionSingh et al. (2010). The design of reward functions should optimize intrinsic motivations (IM), unlike previous literature about IM that focuses on incentivizing exploration which is dangerous in our offline setting, in our paper, we form it as a threefold drive: the drive to restore homeostasis(survival), the drive to lower uncertainty(confidence), and the drive to minimize cost(competence). In Hull’s Drive Reduction TheoryHull (1945), organisms are motivated to maintain physiological equilibrium, we treat homeostasis as a proxy of survival to optimize in the reward function. Unlike curiosity-driven frameworks that encourage exploration under uncertaintyBerlyne (1960); Schmidhuber (1991), in order to design a reward function specific to our offline environmentSingh et al. (2009), we conversely suppress uncertainty due to the sensitivity in critical careDjulbegovic (2011). White (1959) furthers the learning process of effective interaction with the environment, which is transferrable as effective treatment under clinical setting.
@@ 3.2 Potential Based Reward Function Generation
We design a health potential that reflects the quality of physiological states. Past reward functions often define specific rules that directly scores the action, for example, Reward +1+1 if $70<mbp<80$, which will likely cause reward hacking and encouraging actions that can directly increase blood pressure. Instead, we form it as the gradient of potential to capture the patient state change. To bridge the gap between sparse outcome labels and dense clinical supervision, the reward function $R$ is composed of three distinct drives—Survival, Confidence, and Competence leveraging the intrinsic motivation theory Singh et al. (2009). We define the reward function as follows:
$R(s,a,t,s^{\prime},a^{\prime},t^{\prime})=\left[\gamma\Phi(s_{t+1},t+1)-\Phi(s_{t},t)\right]-\lambda\mathcal{C}(a_{t})$ (1)
$\Phi(s_{t},t)=\delta(t)\cdot\sum_{f\in\mathcal{F}}\omega_{f}\cdot\mathcal{S}_{f}(v_{t,f})\cdot\mathcal{U}_{f}(\Delta t_{t,f})$ (2)
@@
We use physiological stabilization as the proxy for survival. By embedding the metadata of critical features as part of the prompt, we personalize the reward design to handle critical patients with comorbidities. The LLM is prompted to design a potential function that rewards states closer to homeostasis. In the offline clinical setting, data irregularity poses a significant risk: a normal heart rate recorded 12 hours ago is not a reliable indicator of current stability. Since we are unable to refresh the data by ordering new lab tests, we utilize the elapsed time $\Delta t$ as indicators of freshness of the data $s_{t}=(o_{t},\Delta t_{t})$. The LLM generates confidence weights for each critical feature that inversely scales the reward based on $\Delta_{t,f}$. This mechanism penalizes stale states, effectively discounting the survival reward when data is uncertain to reduce epistemic uncertainty. We also add a strategic decay to prevent unnecessarily longer ICU stay. To prevent the agent from achieving survival via excessive or harmful interventions, we incorporate a competence component focused on treatment efficiency. Mathematically, this serves as a regularization term: if two treatments result in the same physiological transition $s_{t}\rightarrow s_{t+1}$, the potential function assigns a higher value to the less invasive action, penalizing over-treatment or flailing behavior. By decoupling the accumulating action cost $\mathcal{C}(a)$ from the telescoping potential $\Phi(s,t)$1, we effectively formulate the RL objective as the Lagrangian of a Constrained Markov Decision Process (CMDP). The agent maximizes the Lagrangian $\mathcal{L}=\text{Physiology}-\lambda\cdot\text{Toxicity}$, ensuring that high-stability states are only pursued if the cost of the intervention path does not exceed the value of the stabilization2. Instead of manual heuristic tuning, the LLM acts as the logic designer. We feed the definitions of these three drives along with the feature metadata into the LLM, which autonomously synthesizes the logic into executable Python code for the reward function. To mitigate the risk of hallucination and capture diverse clinical perspectives, we generate $N$ distinct candidate functions in parallel, creating a diverse pool of candidates for the subsequent verification phase.
@@ 3.3 Tri-Drive Reward Selection
Since online verification is ethically prohibitive in clinical settings, we cannot evaluate reward candidates by deploying them on patients. Instead, we propose a set of offline fitness metrics to rigorously assess the quality of each generated reward function $R$. We formulate reward design as a Tri-Drive Selection Problem aiming to balance three conflicting objectives based on trajectory-level labeled ground truth: Survival, Confidence, and Competence Fitness.Survival Fitness ($J_{surv}$) validates whether the induced reward signal correlates with actual patient outcomes. We define a ground-truth trajectory score $G(\tau)$ that consists of both mortality and Sequential Organ Failure Assessment(SOFA) scoreLambden et al. (2019) increase compared to baseline(time when admitted to ICU). The fitness is the correlation $\rho$ between the cumulative reward and $G(\tau)$:
@@
$h(f)=\begin{cases}\mathbf{1}_{f\in\mathcal{I}}+\mathbf{1}_{f\notin\mathcal{I}}\cdot\sigma\left(k\left[0.5-\frac{d(f,\mathcal{I})}{\text{IQR}}\right]\right)&\text{if }T=N,\\ \sigma\left(k\left[0.5-f\right]\right)&\text{if }T=D_{l},\\ \sigma\left(-k\left[0.5-f\right]\right)&\text{if }T=D_{h},\end{cases}$ (7)
@@
where $\sigma(z)=(1+e^{-z})^{-1}$ is the standard sigmoid function, $d(f,\mathcal{I})$ is the distance from feature $f$ to the interval $\mathcal{I}$, IQR is the interquartile range width used for normalization. $\alpha$ and $k$ are cost parameters, $\bar{a}_{t}$ is the magnitude of the intervention. We employ the Non-Dominated Sorting Genetic Algorithm II (NSGA-II) Deb et al. (2002) to rank the generated candidates based on these three objectives. Instead of selecting a single weighted average, we identify the Pareto Frontier where no objective can be improved without degrading another. From this frontier, the solution closest to the utopia point representing the best trade-off.
@@ 4.1 Tasks
Sepsis Treatment We utilize the Medical Information Mart for Intensive Care (MIMIC)-IV v3.1Johnson et al. (2020) database to learn optimal Intravenous (IV) fluid and Vasopressor administration strategies for sepsis patients. We select adult patients (age $\geq$ 18) satisfying the Sepsis-3 criteria Singer et al. (2016), and exclude patients with missing mortality, demographics, and interventions, extreme missingness (>>60% missing data) or short hospital stays (<<24 hours). We define a septic shock starting from 24 hours before diagnosis and ending at 48 hours post-diagnosis. After filtration, 7,501 patients with 46 features are retained in the dataset. The action space is discrete consisting of two primary interventions: IV fluids and Vasopressors. We discretize IV fluid volume and vasopressor dosage into 5 bins each, resulting in a combined action space of $5\times 5=25$ discrete actions.
@@
Mechanical Ventilation We use the multi-center eICU Collaborative Research Database (eICU-CRD)Pollard et al. (2018) to optimize ventilator settings for patients with respiratory failure. The cohort includes adult patients who underwent invasive mechanical ventilation (MV) for at least 24 hours. We select patients aged $\geq$16 who received MV for at least 24 hours. We also exclude patients with missing mortality outcomes, demographics, interventions, and those ventilated for >>14 days. The final cohort contains 2168 patients with 41 features. The action space is discrete, controlling three key ventilator parameters: Positive End-Expiratory Pressure (PEEP), Fraction of Inspired Oxygen (FiO2), and Tidal Volume. We define a discrete action set based on clinically relevant adjustments for each parameter, resulting in a composite discrete action space of 18 actions (PEEP: 2 levels, FiO2: 3 levels, Tidal Volume: 3 levels).
@@
Renal Replacement Therapy We apply our method to the AmsterdamUMCdb Thoral et al. (2021) to optimize continuous renal replacement therapy (RRT) dosing. We select adult patients (age $\geq$ 18) who received CRRT during the ICU stay where the stay is no longer than 60 days. There are 857 patients with 18 features in the cohort. Unlike the previous tasks, the action space here is continuous. The agent controls the effluent flow rate clipped to the range $[0,60]$ ml/kg/h, a critical parameter determining the intensity of dialysis.
@@
For all tasks, clinical time-series data (vitals, labs, medications) are aggregated into 1-hour time steps. Missing values are imputed using forward-filling for time-varying features. For CRRT effluent rate, we imputed values based on RRT sessions, defining a new session whenever the gap between two dialysis records exceeded 8 hours. All continuous state features are normalized to zero mean and unit variance to ensure training stability. For policy training, we use common safe offline RL methods, Batch-Constrained Q-Learning (BCQ)Fujimoto et al. (2019) for discrete action space tasks (Sepsis and Ventilation), and Implicit Q-Learning (IQL)Kostrikov et al. (2021) for continuous action space task (RRT).
@@ 4.2.1 Heuristic Reward Design
We implement three standard reward structures commonly used in clinical RL literature. These serve as the "human-expert" baselines, representing varying degrees of clinical foresight:
Outcome Reward Model (ORM): A purely outcome-oriented function that assigns a non-zero reward only at the terminal step, +100+100 for survival (discharge) and 100-100 for mortality. The value of $\pm 100$ is used in both SepsisKomorowski et al. (2018) and VentilationPeine et al. (2021) tasks in previous literature. Since there is no past work that utilized outcome reward for RRT task, we adopt the same setting of $\pm 100$ for RRT. This baseline tests the agent’s ability to solve the credit assignment problem without intermediate guidance. We explicitly utilized ICU-discharge mortality rather than 90-day mortality as our terminal state definition in the baselines. We argue that an RL agent can only be held accountable for outcomes within its scope of control. An agent penalized for a death occurring 60 days post-discharge may erroneously learn that stabilizing and discharging a patient is suboptimal, leading to conservative strategies.
Process Reward Model (PRM): A process-oriented function that provides immediate feedback based on step-wise improvements in physiological stability. It incentivizes short-term stabilization without long-term survival awareness. We use the reward function fully based on SOFA scoreZhang et al. (2024c) for Sepsis. For Ventilation, though we note that there is one process reward based on ventilator free daysYousuf et al. (2025), the reward calculated on our cohort is not dense enough for efficient policy training, we therefore adopt the intermediate reward from the OPRMLiu et al. (2024). This further indicates that a fixed reward function may not be transferrable between different cohorts even with the same dataset. For RRT, we use the reward function based on Estimated Glomerular Filtration Rate (eGFR)Zhang et al. (2024a).
Outcome Process Reward Model(OPRM): A weighted combination of the sparse and dense signals. This represents the state-of-the-art manual design, aiming to balance immediate physiological regularization with the ultimate goal of patient survival. We select from current SOTA methods Perera et al. (2026)for Sepsis and Liu et al. (2024)for Ventilation. Due to the lack of work on using OPRM for RRT, we choose to combine the PRM(eGFR) with ORM($\pm 100$) as OPRM.